• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ulmer

    Log entry time 15:59:43 on May12,2003

    Entry number 101454

    keyword=Day Shift Summary

    08:30 Septum LHe level OK. Set septum current and correction coils.
    08:50 Ask for beam while septum coming up.
    09:10 Septum is ready at set current.
    09:15 Achieved centered beam at 3 uA, start acquisition on multifoil / Kin 2.11
    09:20 Hall C target warmup starting.
    09:23 Start real acquisition (after adj prescales) - 12C multi
    09:40 Start 3He acq - 12M events above bkgd --> 30M total events --> 2hrs at 3 uA
    with PS1=12.
    10:10 Stop acquisition for beam studies (beam is supposed to be returned at or before 2PM).
    14:00 MCC calls, ready to deliver beam.
    14:08 Beam back, 3 uA and centered, raster at 2.5x3 (on spot++).
    14:19 Start acquisition on 3He, Kin 2.11
    15:16 Start vertical beam scan for Compton (at most 1 mm shift of beam on our target). Stop acquisition and start another run (run 1938).
    15:42 Compton scan finished - start run 1939
    15:52 MCC called: Hall C 15 K target cooldown will commence
    15:56 Stop acquisition. Adjust HV on PMT#3 Gas Cher-Right Arm


    Run Summary:

    Run Tgt Events (All for Kin 2.11, 0 deg, beam 1/2-wave IN)
    ---- ---- -------

    1930 12C
    1931 3He 7.5M
    1937 3He 14M
    1938 3He 5.2M (Compton beam scan during this run: <1 mm shift at tgt)
    1939 3He 3.4M

    Total = 30.1M 3He

    Goal: 30M events (12M above bkgd)

    We're done and now next shift can run with beam 1/2-wave OUT.

    But, first, Bogdan is adjusting HV on PMT#3 gas Cherenkov-Right (after run 1939 ended).